# Conceptualization of this study, Methodology, Software

### v1

@@ 4.4 Multimodal Training
We extend the class-incremental continual learning to a more realistic semi-supervised continual learning (SSCL) setting, where data distributions reflect existing object class correlations between, and among, the labeled and unlabeled data distributions. For our experiment we use 70% labeled data and 30% unlabeled data for the two protocols (random and unseen) discussed above. At task n, we denote batches of labeled data as $X_{n}={(x_{b},y_{b}):b_{i}\in(1,...,b)|y_{b}\in\tau_{n}}$ and batches of unlabeled data as $X_{n}^{{}^{\prime}}={(x_{b}^{{}^{\prime}}):b_{i}\in(1,...,b)}$, where $b$ refers to the batch size. The goal in task $n$ is to learn a model which predicts object class labels for any query input over all classes seen in the current and previous tasks ($\tau_{0}\cup\tau_{1}\cup...\cup\tau_{n}$). During each training step the model receives different classes incrementally over all the experiences for each domain as shown in Figure 1. The two domains contain data from different modalities specifically domain 1 contains the sensor data whereas domain 2 contains the vision data. The data for each task goes through a feature extractor to derive the feature maps from the different layers of the network. The feature maps for both domains are standardized to a fixed dimension of 384 and normalized to fall within the 0-1 range before being fed into the subsequent stages of the algorithm. These feature maps are then used to incrementally train the exFecam algorithm. During training for each task in the second domain, a small subset of the feature maps of the object is stored in a temporary memory buffer to facilitate pseudo-labeling for semi-supervised learning, as previously discussed.
The feature maps are used to compute the mean and covariance matrix for each batch during training. The model performs covariance matrix approximation, shrinkage, and normalization to stabilize the training process Goswami et al.. During the test phase, the extracted feature maps from the test dataset are used to calculate the Mahalanobis distance for all seen classes, and the class with the minimum distance value is selected as the output class. More details about the implementation of the algorithm is provided in the pseudo code. Additionally, we also perform a real-time test for object classification based on the multimodal data using the ROS framework as discussed in section 6 of the paper.
Figure 4: Intra layer feature representation for the custom dataset and the Core50 dataset, we compare the improvement in the accuracy of the FeCAM algorithm on addition of knowledge from the last as well the penultimate layers of the pre-trained feature extractors. The improvement in case of Core50 dataset is considerably more than the custom dataset due to the presence of more diverse information in case of image data.
We further analyzed the intra-layer feature representation for both the custom dataset and the Core50 dataset by comparing the accuracy improvement of the exFeCAM algorithm when incorporating knowledge from the last and penultimate layers of pre-trained feature extractors. The Core50 dataset showed a significantly greater improvement in accuracy compared to the custom dataset. The difference is attributed to the more diverse information contained in the image data of the Core50 dataset.
Figure 5: On the left: radar plot shows the accuracy of the two domains after incremental training on different tasks. The FeCAM algorithm learns the first domain in a class-incremental approach, while it learns the second domain in a combined domain-incremental and class-incremental manner. On the right: the plot shows domain 1 accuracy (in orange) with random gaussian noise added to it, the solid line in grey shows the percentage improvement in the overall accuracy of the experiences after addition of the second domain. The black lines show the error bars for the two cases.
@@ 5 Results
To assess the performance of our algorithm and the effectiveness of the proposed modifications, we evaluate the exFeCAM algorithm on a custom multimodal dataset and the standard Core50 dataset used in the literature. The Core50 dataset was chosen for its non-iid nature and its compatibility for self-supervised learning. We do not explicitly compare our algorithm with other state-of-the-art (SOTA) benchmarks because the focus of this paper is not to establish a new level of accuracy for object classification, but rather to demonstrate the effectiveness of CL in a real world multimodal training setting. Nevertheless, our results are based on improvements to the original FeCAM algorithm, indirectly providing a comparison. All experiments are repeated three times, and we plot the mean and standard deviation of the respective metrics.
Figure 6: Accuracy of the FeCAM algorithm on SSL tasks, the blue and green plot shows the SSL condition for unseen object from the same parent class, with blue indicating SSL enabled and green indicating SSL disabled case. The maroon and red plots represent the SSL condition for randomly selected images of the object, with maroon indicating SSL enabled and red indicating SSL disabled case. Top plot shows the accuracy for the custom multimodal dataset and the botton plot shows the accuracy for the Core50 dataset. The black dotted lines denotes the average accuarcy over all the experiences for different SSL cases.
[…]
@@
Figure 4 (bottom) presents similar information for the Core50 dataset, where we employed the same experimental protocols. The key difference lies in the pre-trained model and the pre-training dataset used in this experiment. However, the improvement in accuracy for the Core50 dataset is minimal compared to the custom dataset. This could be attributed to the significant variations in the objects within each class, making the simple cosine similarity method less effective in this context. A more complex SSL method might be more beneficial for this scenario, which we plan to explore in future experiments.
@@ 6 Real-Time Evaluation With ROS
Robot Operating System (ROS) is a comprehensive framework that comprises a suite of tools, libraries, and protocols aimed at facilitating the development of various robotic systems. The system manages the creation and control of communication between a robot’s peripheral modules, such as sensors, cameras, and actuators, thereby enabling the seamless integration of these components with parallization capabilties.
Figure 8: Real-time images of the object from the camera’s perspective, along with the sensor data from four sensors attached to the soft finger, upon its actuation. These images and sensor readings are directly fed into the FeCAM model after resizing them to the desired dimension. For a single 15-second run, there are a total of 150 sensor signals and 300 images, representing the number of raw data points and corresponding images of the object.
@@ 6.1 ROS Setup
ROS is primarily designed to function optimally on Ubuntu, while only partially compatible with other operating systems, such as Windows and Mac. For our case we use ROS on Ubuntu 20.04 on a Intel Core i7 7th gen CPU @@@@ 2.80 GHz equipped with Nvidia GTX 1080 GPU. We use the Noetic Ninjemys distribution of ROS 1 Koubaa et al. in our experiment. A detailed explanation about the workflow of the project is provided in the flowchart in Figure 7.
Figure 9: Square wave sent to the control box to actuate the soft pneumatic finger. We pressurise the finger for 3 seconds with a maximum pressure of 1.3 bars and then depressurise it for 2 seconds, the cycle is repeated "N" times till we reach the 15 seconds threshold for each run.
The first publisher (P1) publishes the sensor signals coming from the four sensors attached to the soft finger, the second publisher (P2) publishes the video coming from the camera and the third publisher (P3) publishes the command pressure to the control box which inturn actuates the finger. The control frequency of P1 and P2 are 10hz and 20hz, whereas the actuation frequency of P3 is 0.2hz. We further observed that actuating the finger at a more higher frequency constrained it from properly holding the object and thus effected the quality of the generated data from the sensors. The data from P1 and P2 goes to the subscriber which then pre-process the data, extracts the frames from the video in case of P2 and sends it to the respective feature extractors. The output of the feature extractor is a 384-dimensional feature map (in case the intra-layer feature representation is turned on) which goes as an input to the exFecam model for further processing.
@@ 6.2 Real-Time Experiment
To assess the performance of our algorithm, we conduct a quasi real-time experiment using a soft pneumatic gripper equipped with two flex sensors and two force sensors. During the evaluation phase, data from the sensors and the camera are collected in real-time for approximately 15 seconds. These data are treated as two distinct domains and are sent through the pipeline in parallel for further processing and subsequently to the exFecam model for predictions. To generate the predictions we utilize the stored mean and the co-variance matrix from a prior offline incremental experiment conducted for the two domains. Impressively, the mean and covariance matrix occupy only 22.5 MB of memory, making it quite efficient compared to other continual learning strategies
However, since the model was initially trained on images of the objects in a different frame of reference, where the operator extends his arm and smoothly rotates the object in front of the camera. In contrast, during the real-time experiment, the objects are positioned near the soft finger, and a camera captures images from a top-down view. This difference sometimes leads to incorrect predictions by the model. To address this issue, for some objects, we occasionally need to refine the model’s knowledge by conducting a brief training session (lasting around 82 seconds) specifically for that object. Figure 8 shows a snapshot of the real-time data collection from both the sensors and the camera’s perspective. The image shows the progression of the sensor signal and the camera image upon actuating the finger. For each cycle the sensors collect 150 raw values, which are transformed into a single 600-dimensional vector, whereas the camera collects around 300 images of the object from various orientations within the 15-seconds interval.
As discussed above one of the major advantages of this domain incremental training is for autonomous agents in situations where the field of view of one of the sensors gets restricted due to the presence of some obstacle from the environment, in that case the presence of other sensor can compensate for the missing knowledge.
@@ 6.3 Real World Application
A real world application for such a technique can be in companion robots Broadbent et al., where the robot learns new tasks incrementally by interacting with its environment and recognizing different objects using information from multiple sensors available to it. The use of multiple sensors is particularly advantageous in scenarios where a single sensor cannot provide complete information about an object. For instance, a vision camera might struggle to differentiate between two objects with similar structures but with different textures or materials, with one being much softer than the other. In such cases, information from a tactile sensor can be extremely beneficial. Our proposed framework simplifies the addition of new sensors by treating each as a new domain, eliminating the need to retrain the entire algorithm from scratch each time. All the sensors and the algorithm can be synchronized using a mobile version of ROS for embedded systems Min et al., providing a comprehensive solution.
@@ 7 Conclusion and Discussion
In this paper, we present a novel continual learning method that combines the informations from different sensors in a domain incremental fashion while also learning new classes or objects in a class incremental manner. We enhanced the original FeCAM algorithm by incorporating online batch-wise training, intra-layer feature representation, and SSL capabilities thereby making the algorithm suitable for our scenario. The exFeCAM algorithm employs cosine similarity based feature matching to perform pseudo labeling when encountering unlabeled data. We validated the algorithm on two different modalities of data namely sensor signal based data and image based data, and demonstrated the ability of the algorithm to learn the two modalities incrementally. We also conducted an ablation study to show the contribution of different parts of the algorithm. Additionally, we performed a real-time experiment on a soft pneumatic gripper equipped with four sensors and an external camera system, all synchronised using the ROS framework. As part of our future work, we aim to enhance the exFeCAM algorithm further by adding forward and backward knowledge capabilities, either through interactions between class prototypes or by training a simple multilayer perceptron (MLP) network to learn these interactions. Moreover, currently we need to manually select the prediction domain modality in the presence of noise; in the future, we aim to enable the algorithm to automatically determine this by calculating the confidence of each prediction from the network.
CRediT authorship contribution statement
Nilay Kushawaha: Writing - original draft, Methodology, Investigation, Experimentation, Formal Analysis, Conceptualization. Egidio Falotico: Supervision, Funding acquisition, Project administration.
Declaration of competing interest
@@
Data availability
Data will be made available on request. The code is available on the github repo : https://github.com/nilay121/Multimodal-Fusion-using-CL
Acknowledgment
We acknowledge the contribution from the Italian National Recovery and Resilience Plan (NRRP), M4C2, funded by the European Union–NextGenerationEU (Project IR0000011, CUP B51E22000150006, "EBRAINS-Italy")

### v2

@@ 4.4 Multimodal Training
We extend the traditional class-incremental CL framework to a more practical and challenging semi-supervised continual learning (SSCL) scenario. In this setting, the data distributions are designed to mirror real-world correlations among object classes by incorporating both labeled and unlabeled samples. For our experiments, we use a data split consisting of 70% labeled and 30% unlabeled data, applied under two protocols (random and unseen) as described in the previous section. At each task n, we represent the batches of labeled data as
$X_{n}=\{(x_{b},y_{b}):b_{i}\in(1,...,b),\,y_{b}\in\tau_{n}\}$
and the batches of unlabeled data as
$X_{n}^{{}^{\prime}}=\{x_{b}^{{}^{\prime}}:b_{i}\in(1,...,b)\},$
where $b$ denotes the batch size and $\tau_{n}$ indicates the set of classes associated with task $n$. The main goal during each task is to train a model that can accurately classify any query instance drawn from the union of all classes encountered in the current and preceding tasks, i.e., $\tau_{0}\cup\tau_{1}\cup\dots\cup\tau_{n}$. As training progresses, the model encounters new classes incrementally within each domain, as illustrated in Figure 1, where each domain corresponds to a different data modality. For each task, input data are passed through a feature extractor that produces feature maps from various layers of the network. These feature maps are standardized to a fixed dimension and normalized to 0–1 range before being fed into subsequent stages of the pipeline. These processed feature representations are then used to incrementally train the exFeCAM algorithm. To support the semi-supervised aspect of learning, a small subset of feature maps corresponding to the objects in each task is stored temporarily in a memory buffer. This buffer is utilized for pseudo-labeling of unlabeled data, as discussed in earlier sections.
During the training phase, the algorithm computes the mean and covariance matrix for each batch of data and apply various techniques such as covariance matrix approximation, shrinkage, and normalization to ensure robust and stable learning, as described in the paper [8]. In the testing phase, feature maps for the test instances are extracted using the appropriate feature extractor, and classification is performed by computing the Mahalanobis distance between each test sample and the class prototypes learned during training. The class corresponding to the minimum distance is selected as the predicted label. Further details about the algorithm implementation are provided in the corresponding pseudo-code.
Figure 5: On the left, the radar plot illustrates the accuracy of various algorithms after incremental training on each experience for the custom multimodal dataset. The accuracy per experience represents the average performance across two domains: visual and sensor signals. On the right, the radar plot presents the per-experience accuracy for the VGGSound dataset, where the two domains consist of visual and audio data.
@@ 5 Results
To evaluate the performance of our algorithm and the effectiveness of the proposed modifications, we assess the exFeCAM algorithm on both a custom multimodal dataset and the VGGSound dataset. It is important to note that the primary focus of this paper is not to establish a new level of accuracy for object classification, but rather to demonstrate the effectiveness of CL in a challenging multimodal training setting. All experiments are conducted three times, and we report the mean and standard deviation of the relevant metrics.
Table 2 compares the accuracy of various algorithms across the two multimodal datasets. Our proposed method achieves a significant improvement over all benchmark algorithms, which can be attributed to the integration of SSL and intra-layer feature representation (ILFR) capabilities in our architecture. Additionally, the vanilla FeCAM algorithm produces competitive results, likely due to it’s offline training mechanism. In contrast, our algorithm employs batch-wise online training, making it more relevant for real-world scenarios. FeTrIL algorithm also achieves a comparable accuracy, mostly attributed due to the presence of the pseudo-rehearsal technique to compensate forgetting of previous knowledge.
Figure 4 (left) illustrates the enhancement in our algorithm’s prediction accuracy with the integration of intra-layer feature representation (ILFR). The accuracy is measured after the model has been incrementally trained on all tasks. For the custom dataset (represented in blue), the first five experiences depict the percentage improvement in domain 1 data (vision), while the subsequent five experiences reflect the accuracy gains in domain 2 (tactile sensor). Each experience consists of two distinct classes for both domains. On average, incorporating ILFR results in an overall accuracy improvement of approximately 3.12 $\pm$ 1.09 compared to the baseline without ILFR. A similar trend is observed in the VGGSound dataset (also shown in blue), where the initial five experiences correspond to image data, followed by five experiences with the associated audio data. In this case, the overall accuracy improvement is approximately 4.90 $\pm$ 1.20, further validating the benefits of ILFR.
Figure 4 (right) demonstrates the advantage of training on a new domain when the information from the original domain (signals from the tactile sensor) is noisy. To simulate this scenario, gaussian noise is added to the sensor signal data. The solid orange line indicates the accuracy across different experiences for the signal domain with noisy sensor data, while the dashed orange line shows the original accuracy without noise. The solid black line indicates the improvement in accuracy when image data is incrementally introduced to mitigate the performance degradation caused by noisy sensor signals.
The CL capabilities of exFeCAM, along with the benchmark algorithms, are presented in Figure 5. The spider plot depicts the combined accuracy for both domains, demonstrating that our proposed algorithm consistently outperforms all other methods across all experiences. The naive-concat and naive-add benchmark methods perform well in the initial experience but suffer significant accuracy degradation as the number of classes increases due to catastrophic forgetting. The joint method, which trains on all data at once, serves as an upper bound on per-experience accuracy. The FeTrIL algorithm initially shows results comparable to vanilla FeCAM, but it’s performance declines in later tasks, likely due to the moderate quality of the pseudo-rehearsal samples generated during training on new tasks.
Figure 6: Accuracy of the exFeCAM algorithm across two different SSL scenarios. The blue plots correspond to the custom dataset, while the green plots represent the VGGSound dataset. Bar plots with solid lines indicate the condition where SSL is integrated into the algorithm, whereas dotted lines represent the condition without SSL.
Additionally, we conducted a separate SSL experiment exclusively on the visual modality from both datasets, implementing two distinct types of experiments within the SSL protocol, as described in Section 3.3. Figure 6 presents the accuracy results for the custom dataset and the VGGSound dataset under both scenarios. The blue bar plot with a solid boundary represents the accuracy of the exFeCAM algorithm for the unseen object and random object scenarios with SSL enabled, while the dotted-boundary plot indicates the accuracy for the same scenarios without SSL. Similarly, the accuracy results for the VGGSound dataset are depicted using green bar plots. However, the overall accuracy of the VGGSound dataset is lower compared to the custom dataset, possibly due to the increased complexity and diversity of backgrounds in the extracted frames, whereas the custom dataset primarily consists of images with a relatively uniform background.
@@ 6 Real-Time Evaluation With ROS
Robot Operating System (ROS) [9] is a comprehensive framework that comprises a suite of tools, libraries, and protocols aimed at facilitating the development of various robotic systems. The system manages the creation and control of communication between a robot’s peripheral modules, such as sensors, cameras, and actuators, thereby enabling the seamless integration of these components with parallization capabilties.
Figure 8: Real-time images of the object from the camera’s perspective, along with the sensor data from four sensors attached to the soft finger, upon it’s actuation. These images and sensor readings are directly fed into the FeCAM model after resizing them to the desired dimension. For a single 15-second run, there are a total of 150 sensor signals and 300 images, representing the number of raw data points and corresponding images of the object.
@@ 6.1 ROS Setup
ROS is primarily designed to function optimally on Ubuntu, while only partially compatible with other operating systems. For our case we use ROS on Ubuntu 20.04 on a Intel Core i7 7th gen CPU @@ 2.80 GHz equipped with Nvidia GTX 1080 GPU. We use the Noetic Ninjemys distribution of ROS 1 [9] in our experiment. A detailed explanation about the workflow of the project is provided in the flowchart in Figure 7.
Figure 9: Real-time object classification accuracy of the two modalities for four different objects. The grey plot represents the prediction accuracy in the image domain, while the pink plot highlights the accuracy for the sensor domain.
The first publisher (P1) publishes the sensor signals coming from the four sensors attached to the soft finger, the second publisher (P2) publishes the video coming from the camera and the third publisher (P3) publishes the command pressure to the control box which inturn actuates the finger. The control frequency of P1 and P2 are 10hz and 20hz, whereas the actuation frequency of P3 is 0.2hz. We further observed that actuating the finger at a more higher frequency constrained it from properly holding the object and thus effected the quality of the generated data from the sensors. The data from P1 and P2 goes to the subscriber which then pre-process the data, extracts the frames from the video in case of P2 and sends it to the respective feature extractors. The output of the feature extractor is a 384-dimensional feature map for both the domains which then goes as an input to the exFecam model for further processing and prediction.
@@ 6.2 Real-Time Experiment
To assess the performance of our algorithm, we conduct a quasi real-time experiment using a soft pneumatic gripper equipped with two flex sensors and two force sensors. During the evaluation phase, data from the tactile sensors and the camera are collected in real-time for approximately 15 seconds. These multimodal data are treated as two distinct domains and are sent through the dedicated pipeline in parallel for further processing and subsequently to the exFecam model for predictions as shown in Figure 7. To generate the predictions we utilize the stored mean and the co-variance matrix computed during the training phase. Impressively, the mean and covariance matrix occupy only 22.5 MB of memory, making it quite efficient compared to other CL strategies.
However, since the model was initially trained on images of the objects in a different frame of reference, where the operator extends his arm and smoothly rotates the object in front of the camera. In contrast, during the real-time experiment, the objects are positioned near the soft finger, and a camera captures images from a top-down view. This difference sometimes leads to incorrect predictions by the model. To address this issue, for some objects, we occasionally need to refine the model’s knowledge by conducting a brief training session (lasting around 82 seconds) specifically for that object. Figure 8 shows a snapshot of the real-time data collection from both the tactile sensors and the camera’s perspective. The image shows the progression of the sensor signal and the camera image upon actuating the finger. For each cycle the sensors collect 150 raw values, which are transformed into a single 600-dimensional vector, whereas the camera collects around 300 images of the object from various orientations within the 15-seconds interval.
As discussed earlier, a significant advantage of domain-incremental CL training is it’s ability to incorporate new data modalities incrementally. This is particularly beneficial in scenarios where the input from one sensor becomes unreliable, either due to obstructions in the environment that limit it’s field of view or due to noise introduced by external factors. This advantage is further illustrated in Figure 9, where we evaluate object classification accuracy in real-time for four different objects: stapler, earpod, bottle, and marker pen. The results reveal that for the ”earpod” and ”marker pen”, the tactile sensor modality achieves notably higher classification accuracy than the corresponding visual modality. This discrepancy is likely caused by occlusions from the operator’s hand during image capture or limitations in the visual feature extractor’s ability to accurately represent small objects. In such cases, the availability of a second modality (here tactile sensor) effectively compensates for the missing or unreliable information, thereby improving overall classification performance. All the experiments were repeated 5 times for each of the four objects in real-time, and the final prediction accuracy was computed accordingly.
@@ 6.3 Real World Application
A real world application for such a technique can be in companion robots [58], where the robot learns new tasks incrementally by interacting with it’s environment and recognizing different objects using information from multiple sensors available to it. The use of multiple sensors is particularly advantageous in scenarios where a single sensor cannot provide complete information about an object. For instance, a vision camera might struggle to differentiate between two objects with similar structures but with different textures or materials, with one being much softer than the other. In such cases, information from a tactile sensor can be extremely beneficial. Our proposed framework simplifies the addition of new modalities by treating each as a new domain, eliminating the need to re-train the entire algorithm from scratch each time.
@@ 7 Conclusion and Discussion
In this paper, we present a method for incrementally learning new data modalities without the need to re-train the model from scratch each time a new modality becomes available. Our approach combines class-incremental CL to progressively learn new object classes and domain-incremental CL to incorporate additional modalities for existing classes. To support this dual-incremental setup, we build upon the original FeCAM algorithm by introducing key enhancements, including online batch-wise training, intra-layer feature representation, and SSL capabilities, resulting in a more robust and versatile framework.
Our method employs cosine similarity based feature matching to generate pseudo-labels for unlabeled data, enabling the model to effectively leverage both labeled and unlabeled examples during training. We evaluate the proposed algorithm on two multimodal datasets of varying complexity and demonstrate it’s ability to incrementally learn new classes in different domains. To better understand the contribution of individual components, we also conduct a comprehensive ablation study. To validate real-world applicability, we perform a real-time experiment using a soft pneumatic gripper equipped with four tactile sensors and an external camera. This experiment highlights the practical advantages of our multimodal CL approach in a realistic setting.
As part of our future work, we plan to further improve the exFeCAM algorithm by incorporating forward and backward knowledge transfer, either through interactions between class prototypes or by training a multilayer perceptron (MLP) network to model these interactions. Additionally, instead of relying on a simple average of predictions from both modalities, we aim to adopt a more refined approach by using a weighted averaging scheme, where a separate network will estimate the confidence of each prediction to determine the appropriate weights.
Acknowledgements
We acknowledge the contribution from the Italian National Recovery and Resilience Plan (NRRP), M4C2, funded by the European Union–NextGenerationEU (Project IR0000011, CUP B51E22000150006, ”EBRAINS-Italy”)
Conflicts of Interest
@@ Acknowledgements
Data Availability Statement
Data will be made available from the corresponding author upon reasonable request.
Ethical Approval Statement
All experiments were conducted solely by the author, and no other human participants were involved in the study. The author therefore gives informed written consent and confirms that no ethics committee approval is required, as per university guidelines.